5,1 is a solution to the equation 2x-3y=7. What does this mean?
Found 2 solutions by jim_thompson5910, stanbon:
Answer by jim_thompson5910(28595)

If you're given the point (5,1), this means that x=5 and y=1. So simply plug in x=5 and y=1 into 2x-3y=7. If you end up with an identity, then that will confirm that (5,1) is a solution. If (5,1) is a solution, then that means that the point (5,1) lies on the line 2x-3y=7 (that's only if it's a solution).


Note: an identity is an equation that is true for ALL values of 'x'. A simple example would be 0=0 (0 is always equal to itself). So if you get some number equal to itself, then you have an identity.

Answer by stanbon(57361)

5,1 is a solution to the equation 2x-3y=7. What does this mean?
-------------------------
It means that if you you replace x by 5 and replace y by 1
you get a true statement.
----
Try it:
2*5 - 3*1 = 7
10-3 = 7
7 = 7
===============
Cheers,
stan H.